A compound has the empirical formula of C2H3O and a molar mass of 172g. What is it’s molecular formula


1
Expert's answer
2021-03-22T06:39:43-0400

Empirical formula of the compound: "C_2H_3O"


Suppose, its molecular formula is "(C_2H_3O" )n.

So, we can write the following equation to find out the value of n.


"n\times(2×12+3×1+16) = 172"

Or, 43n = 172


Or, n = "\dfrac{172}{43}" =4


So, the molecular formula of the compound is "C_8H_{12}O_4"